# Explaining Similarity in Vision–Language Encoders with Weighted Banzhaf Interactions

### v1

@@ Abstract
Language–image pre-training (LIP) enables the development of vision–language models capable of zero-shot classification, localization, multimodal retrieval, and semantic understanding. Various explanation methods have been proposed to visualize the importance of input image–text pairs on the model’s similarity outputs. However, popular saliency maps are limited by capturing only first-order attributions, overlooking the complex cross-modal interactions intrinsic to such encoders. We introduce faithful interaction explanations of LIP models (FIxLIP) as a unified approach to decomposing the similarity in vision–language encoders. FIxLIP is rooted in game theory, where we analyze how using the weighted Banzhaf interaction index offers greater flexibility and improves computational efficiency over the Shapley interaction quantification framework. From a practical perspective, we propose how to naturally extend explanation evaluation metrics, like the pointing game and area between the insertion/deletion curves, to second-order interaction explanations. Experiments on MS COCO and ImageNet-1k benchmarks validate that second-order methods like FIxLIP outperform first-order attribution methods. Beyond delivering high-quality explanations, we demonstrate the utility of FIxLIP in comparing different models like CLIP vs. SigLIP-2 and ViT-B/32 vs. ViT-L/16.
@@ 1 Introduction
Contrastive language–image pre-training (CLIP, Radford et al., 2021) revolutionized computer vision by learning data representations well-performing in a plethora of downstream tasks. Scaling the training to predict similarity between image and text, combined with continuous architectural improvements (e.g. SigLIP, Zhai et al., 2023) and increasing the size of datasets, leads to the development of powerful vision–language encoders (VLEs, Zhai et al., 2023; Wan et al., 2024; Chun et al., 2025; Tschannen et al., 2025). These rather opaque encoders become the backbone components in large vision–language models capable of actual conversation and reasoning (Alayrac et al., 2022; Li et al., 2023; Liu et al., 2023), and are increasingly applied in high-stakes decision-making, like in the case of medical imaging (Ikezogwo et al., 2023; Zhang et al., 2024; Zhao et al., 2025).
@@
The limitations in the way CLIP represents the visual world have been extensively studied (Yuksekgonul et al., 2023; Lewis et al., 2024; Udandarao et al., 2024; Covert et al., 2025; Xiao et al., 2025). For example, a recent study by Tong et al. (2024) suggests vision–language models are limited by the systematic shortcomings of CLIP, e.g. they fail when questioned about directions, quantity of objects in an image, or recognizing the presented text in a visual form. The emergence of “CLIP-blind pairs”, inputs that CLIP perceives as similar despite their clear differences, could lead to critical errors when applied in medical visual question answering and treatment recommendation (Zhao et al., 2025). One way of validating similarity predictions at inference time is explaining the encoders with saliency maps (Morch et al., 1995; Chefer et al., 2021; Wang et al., 2023; Zhao et al., 2024; Li et al., 2025). However, these first-order importance scores are limited by their unimodal view of the model’s output (Agarwal, 2025). Vision–language encoders cannot be faithfully explained without taking into account the cross-modal interactions between image–text pairs (Liang et al., 2023; Joukovsky et al., 2023).
@@
We thus propose a model-agnostic approach to interpreting VLEs based on a game-theoretical perspective where input tokens form coalitions of players in a cooperative game (Lundberg and Lee, 2017; Fumagalli et al., 2023). In doing that, we extend the faithful Banzhaf interaction index (Tsai et al., 2023) to a weighted case (Marichal and Mathonet, 2011), addressing the emerging requirement for controllable sparsity in coalition sampling (Hase et al., 2021; Hama et al., 2023; Bluecher et al., 2024; Zheng et al., 2025). Specifically, sparse inputs with many masked tokens become out-of-distribution, i.e. images become unrecognizable and text captions become ambiguous. Figure 1 shows a high-level view of our approach. We sample uniformly $m^{2}$ variations of the input image–text pair. Each mask is sampled uniformly with probability $p$ for each token in a mask, and we take all pairwise combinations between $m$ masked images and $m$ masked texts. Then, we apply the explained model to predict similarity values for each of the input variations. Finally, we apply weighted least squares to approximate the model’s predictions from sampled binary masks. The resulting explanation assigns attribution scores to each token and interaction scores to each pair of tokens in an image–text pair.
@@ 2 Related Work
Attribution explanations of CLIP. Multimodal explanations have been studied in various tasks including visual question answering (Park et al., 2018; Hessel and Lee, 2020), visual reasoning (Lyu et al., 2022), video classification (Wu and Mooney, 2019), and sentiment analysis (Wang et al., 2022). Our work considers explaining the bi-modal encoder prediction in the means of input attributions, specifically image and text token attributions of similarity predictions in CLIP. In this context, existing gradient-based (Zhao et al., 2024), attention-based (Chefer et al., 2021; Li et al., 2025), and information-based (Wang et al., 2023; Zhu et al., 2025) methods approximate only first-order attributions visualized as saliency maps. Yet, Liang et al. (2023) perform extensive user studies showing that visualizing the second-order attributions (pairwise interactions) is necessary for understanding complex multimodal models. Joukovsky et al. (2023) propose to approximate these interactions with bilinear models (also for image–image encoders, Sammani et al., 2024). We formalize the problem in the means of game theory and propose an efficient approximation of cross-modal and intra-modal interactions (Figure 1), discussing its several desirable theoretical properties. In concurrent work, Möller et al. (2025) propose to approximate only cross-modal interactions based on the Integrated Hessians methodology (Janizek et al., 2021). Our proposed evaluation metrics for interaction-based explanations can be used to empirically cross-compare the above-mentioned attribution explanations.
@@
Mechanistic interpretability of CLIP. Our research is orthogonal to work on the concept-based interpretability of the CLIP’s internal representations. Research in this direction considers explaining particular neurons (Goh et al., 2021), concept-based image classification (Sammani and Deligiannis, 2024), linear probing (Bhalla et al., 2024), and training sparse autoencoders (Lim et al., 2025; Zaigrajew et al., 2025). Gandelsman et al. (2024) investigate how the individual model’s components affect its final representation. Beyond CLIP, Balasubramanian et al. (2024) analyze interpreting image representations via text in alternative vision transformer architectures (e.g. DINO, Caron et al., 2021).
@@
Shapley and Banzhaf interactions in machine learning. We build on the developments using game-theoretic interaction indices to understand complex machine learning models (Sundararajan et al., 2020; Bordt and von Luxburg, 2023; Pelegrina et al., 2023; Fumagalli et al., 2023; Tsai et al., 2023). Our goal in this work is not to compare with the most recent advancements in approximating the interaction index (Kang et al., 2024; Musco and Witter, 2025; Enouen and Liu, 2025), but rather to faithfully explain the similarity predicted by VLEs. From the first-order perspective, Parcalabescu and Frank (2023) apply the Shapley value to measure the importance each modality has on various tasks solved by CLIP. Tsai et al. (2023) and Fumagalli et al. (2024) propose a weighted least squares regression approximation of higher-order interactions, which we incorporate as a baseline in our methodology. Recently, Kang et al. (2025) applied a sparse Fourier transform to scale feature interaction explanations for large language models, while we explain vision–language encoders instead. Related to our contribution of weighted Banzhaf interaction explanations is work on using weighted Banzhaf values to estimate data valuation scores (Li and Yu, 2023, 2024), which are the importance that training data points or subsets have on the model’s performance. For a comprehensive overview of work in this direction, see (Rozemberczki et al., 2022; Muschalik et al., 2024) and references given there.
@@ 3 A Game-theoretic Explanation of Similarity in Vision–Language Encoders
An interaction explanation can be viewed as a complete graph with weighted nodes and edges, similar to the SI-Graph (Muschalik et al., 2025), or a symmetric quadratic matrix of dimension $n_{\mathcal{I}}+n_{\mathcal{T}}$, where the diagonal represents token attributions. To compute explanations, we mask tokens with a pre-defined baseline, depending on the encoder’s architecture, e.g. attention masking for text tokens (we defer implementation details to Appendix B). The masking operator is used to define the FIxLIP game that captures all possible masks.
@@ 3.1 FIxLIP-$\boldsymbol{p}$ explanations via weighted faithful Banzhaf interactions
In this section, we formally define the FIxLIP explanation as an instance of Definition˜1, which faithfully describes the similarity of all masked inputs $\nu(M)$. We thus define $\mathbf{e}$ as an additive explanation with interaction terms that recovers $\nu(M)\approx\hat{\nu}_{\mathbf{e}}(M)$ via a $2$-additive game (Grabisch, 2016) as
@@
The FIxLIP-$p$ explanation is the most faithful approximation of $\nu$ with respect to $\mathfrak{F}_{p}$. Its interactions correspond to the weighted Banzhaf interactions (Marichal and Mathonet, 2011). For $p=0.5$, FIxLIP-$p$ is the faithful Banzhaf interaction index (Tsai et al., 2023) with a known analytical solution (Hammer and Holzman, 1992).
@@
Weighted Banzhaf interactions define a cardinal-probabilistic interaction index that satisfies the linearity, symmetry, and dummy axiom (Fujimoto et al., 2006). Alternatively, faithful Shapley interactions optimize a variant of the faithfulness metric (Tsai et al., 2023). In our context, weighted Banzhaf interactions have two benefits: First and foremost, the hyperparameter $p$ allows for prioritizing masks that are expected to better reflect in-distribution inputs, and provides a flexible explanation framework with intuitive weights. Second, it is unknown if the Shapley weights can be factored into independent distributions for each modality – a crucial requirement for the cross-modal estimator introduced in Section˜3.2.
@@
For $p=0.5$, this estimator reduces to Faith-Banzhaf (Tsai et al., 2023) – a variant of KernelSHAP (Lundberg and Lee, 2017).
@@ 4 Evaluation Metrics for Interaction Explanations of VLE Predictions
In this section, we derive three evaluation metrics for explanations that may include second-order interactions. First, we evaluate the faithfulness of the approximation $\hat{\nu}_{\mathbf{e}}$ to $\nu$, since the basis of explaining VLEs is the masking operator that defines the FIxLIP game $\nu$. Note that conventionally this is done with an $R^{2}$ coefficient (Kang et al., 2024; Muschalik et al., 2024) or mean squared error (Fumagalli et al., 2023; Enouen and Liu, 2025; Musco and Witter, 2025), but we need to rely on rankings to compare attribution methods like saliency maps. Second, we generalize the area between insertion and deletion curves (AID, Hama et al., 2023; Zhao et al., 2024), aka remove least/most important first (Bluecher et al., 2024), a popular faithfulness metric for token attribution methods. Finally, we extend the well-established pointing game evaluation of faithfulness from the image classification literature (Bohle et al., 2021, 2024; Arya et al., 2024; Möller et al., 2025). In the latter two metrics, we need to assume $\mathbf{e}_{\{i,j\}}=0$ for the first-order attribution explanations without interactions.
@@ 5 Experiments
Setup. In experiments, we empirically validate the performance of FIxLIP with three metrics defined in Section 4, measure its computational efficiency, and demonstrate its utility in visual explanation of VLEs. We mainly use the openly available pre-trained CLIP models (Radford et al., 2021) of two sizes: ViT-B/32 with $7\times 7$ image patches and ViT-B/16 with $14\times 14$. Moreover, we demonstrate the broader applicability of FIxLIP to explain SigLIP (Zhai et al., 2023) and SigLIP-2 (Tschannen et al., 2025) up to the ViT-L/16 variant with $16\times 16$ patches. We rely on two openly available datasets commonly used in explainability research: MS COCO (Lin et al., 2014) and ImageNet-1k (Deng et al., 2009); the latter specifically to design the pointing game evaluation considering zero-shot classification.
In quantitative evaluation, we compare FIxLIP to a few representative baselines: the first attribution method for CLIP abbreviated as GAME (Chefer et al., 2021), a state-of-the-art first-order attribution method Grad-ECLIP (Zhao et al., 2024), and a recently released second-order interaction method XCLIP (Möller et al., 2025). For comparisons of Grad-ECLIP and XCLIP with previous baselines, refer to the appropriate benchmarks (Zhao et al., 2024; Möller et al., 2025). Figure 2 conceptually compares FIxLIP to baselines using the image–text input example from Figure 1. Furthermore, we naturally demonstrate the improvement between FIxLIP and first-order Shapley/Banzhaf values. In that, we effectively scale FIxLIP to approximate second-order explanations with a budget of over $10^{6}$ model inferences, which far exceeds related work. For FIxLIP-$p$, we use the cross-modal estimator with a budget of $2^{21}$, whereas FIxLIP with Shapley interactions uses the model-agnostic estimator with a budget of $2^{17}$, yielding approximately similar runtime. We mainly experiment with $p\in\{0.3,0.5,0.7\}$; note that there is no computational overhead for different $p$ settings. Further details for the experimental setup are provided in Appendix C.
Figure 2: Visual comparison between FIxLIP and baselines. First-order attribution methods like GAME and Grad-ECLIP lack the tools to faithfully explain complex similarity predictions of vision–language encoders like CLIP. Notably, in this example, the text token antRefer to caption is the most important for the similarity prediction. One of the differences from XCLIP is that we include intra-modal and main effects in the approximation, which are crucial for obtaining faithful interaction explanations.

### v2

@@ Abstract
Language–image pre-training (LIP) enables the development of vision–language models capable of zero-shot classification, localization, multimodal retrieval, and semantic understanding. Various explanation methods have been proposed to visualize the importance of input image–text pairs on the model’s similarity outputs. However, popular saliency maps are limited by capturing only first-order attributions, overlooking the complex cross-modal interactions intrinsic to such encoders. We introduce faithful interaction explanations of LIP models (FIxLIP) as a unified approach to decomposing the similarity in vision–language encoders. FIxLIP is rooted in game theory, where we analyze how using the weighted Banzhaf interaction index offers greater flexibility and improves computational efficiency over the Shapley interaction quantification framework. From a practical perspective, we propose how to naturally extend explanation evaluation metrics, such as the pointing game and area between the insertion/deletion curves, to second-order interaction explanations. Experiments on the MS COCO and ImageNet-1k benchmarks validate that second-order methods, such as FIxLIP, outperform first-order attribution methods. Beyond delivering high-quality explanations, we demonstrate the utility of FIxLIP in comparing different models, e.g. CLIP vs. SigLIP-2.
@@ 1 Introduction
Contrastive language–image pre-training [CLIP, radford2021learning] revolutionized computer vision by learning data representations well-performing in a plethora of downstream tasks. Scaling the training to predict similarity between image and text, combined with continuous architectural improvements [e.g. SigLIP, zhai2023siglip] and increasing the size of datasets, leads to the development of powerful vision–language encoders [VLEs, zhai2023siglip, wan2024locca, chun2025probabilistic, tschannen2025siglip2]. These rather opaque encoders become the backbone components in large vision–language models capable of actual conversation and reasoning [alayrac2022flamingo, li2023blip, liu2023visual], and are increasingly applied in high-stakes decision-making, like in the case of medical imaging [ikezogwo2023quilt1m, zhang2024biomedclip, zhao2025clip].
@@
The limitations in the way CLIP represents the visual world have been extensively studied [yuksekgonul2023when, lewis2024does, udandarao2024visual, covert2025locality, xiao2025flairvlm]. For example, a recent study by tong2024eyes suggests vision–language models are limited by the systematic shortcomings of CLIP, e.g. they fail when questioned about directions, quantity of objects in an image, or recognizing the presented text in a visual form. The emergence of “CLIP-blind pairs”, inputs that CLIP perceives as similar despite their clear differences, could lead to critical errors when applied in medical visual question answering and treatment recommendation [zhao2025clip]. One way of validating similarity predictions at inference time is explaining the encoders with saliency maps [morch1995, chefer2021generic, wang2023visual, zhao2024gradeclip, li2025closer]. However, these first-order importance scores are limited by their unimodal view of the model’s output [agarwal2025rethinking]. Vision–language encoders cannot be faithfully explained without taking into account the cross-modal interactions between image–text pairs [liang2023multiviz, joukovsky2023model].
@@
We thus propose a model-agnostic approach to interpreting VLEs based on a game-theoretical perspective where input tokens form coalitions of players in a cooperative game [lundberg2017unified, fumagalli2023shapiq]. In doing that, we extend the faithful Banzhaf interaction index [tsai2023faith] to a weighted case [marichal2011weighted], addressing the emerging requirement for controllable sparsity in coalition sampling [hase2021outofdistribution, naofumi2023deletion, bluecher2024decoupling, zheng2025ffidelity]. Specifically, sparse inputs with many masked tokens become out-of-distribution, i.e. images become unrecognizable and text captions become ambiguous. Figure 1 shows a high-level view of our approach. We sample uniformly $m^{2}$ variations of the input image–text pair. Each mask is sampled uniformly with probability $p$ for each token in a mask, and we take all pairwise combinations between $m$ masked images and $m$ masked texts. Then, we apply the explained model to predict similarity values for each of the input variations. Finally, we apply weighted least squares to approximate the model’s predictions from sampled binary masks. The resulting explanation assigns attribution scores to each token and interaction scores to each pair of tokens in an image–text pair.
@@ 2 Related Work
Attribution explanations of CLIP. Multimodal explanations have been studied in various tasks, including visual question answering [park2018multimodal, hessel2020multimodal], visual reasoning [lyu2022dime], video classification [wu2019faithful], and sentiment analysis [wang2022m2lens]. Our work considers explaining the bi-modal encoder prediction in the means of input attributions, specifically image and text token attributions of similarity predictions in CLIP. In this context, existing gradient-based [zhao2024gradeclip], attention-based [chefer2021generic, li2025closer], and information-based [wang2023visual, zhu2025narrowing] methods approximate only first-order attributions visualized as saliency maps. Yet, liang2023multiviz perform extensive user studies showing that visualizing the second-order attributions (pairwise interactions) is necessary for understanding complex multimodal models. joukovsky2023model propose to approximate these interactions with bilinear models [also for image–image encoders, sammani2024visualizing]. We formalize the problem in the means of game theory and propose an efficient approximation of cross-modal and intra-modal interactions (Figure 1), discussing its several desirable theoretical properties. In concurrent work, moeller2025explaining propose to approximate only cross-modal interactions based on the Integrated Hessians methodology [janizek2021explaining]. Our proposed evaluation metrics for interaction-based explanations can be used to empirically cross-compare the above-mentioned attribution explanations.
@@
Mechanistic interpretability of CLIP. Our research is orthogonal to work on the concept-based interpretability of CLIP’s internal representations. Research in this direction considers explaining particular neurons [goh2021multimodal], concept-based image classification [sammani2024interpreting], linear probing [bhalla2024interpreting], and training sparse autoencoders [lim2025sparse, zaigrajew2025interpreting]. g2024interpreting investigate how the individual model’s components affect its final representation. Beyond CLIP, balasubramanian2024decomposing analyze interpreting image representations via text in alternative vision transformer architectures [e.g. DINO, caron2021emerging].
@@
Shapley and Banzhaf interactions in machine learning. We build on the developments using game-theoretic interaction indices to understand complex machine learning models [sundararajan2020shapley, bordt2023shapley, pelegrina2023kadditive, fumagalli2023shapiq, tsai2023faith]. Our goal in this work is not to compare with the most recent advancements in approximating the interaction index [kang2024learning, musco2025provably, enouen2025instashap], but rather to faithfully explain the similarity predicted by VLEs. From the first-order perspective, parcalabescu2023mmshap, parcalabescu2025do apply the Shapley value to measure the importance each modality has on various tasks solved by CLIP. tsai2023faith and fumagalli2024kernelshapiq propose a weighted least squares regression approximation of higher-order interactions, which we incorporate as a baseline in our methodology. Recently, kang2025spex applied a sparse Fourier transform to scale feature interaction explanations for large language models, while we explain vision–language encoders instead. Related to our contribution of weighted Banzhaf interaction explanations is work on using weighted Banzhaf values to estimate data valuation scores [li2023robust, li2024one], which are the importance that training data points or subsets have on the model’s performance. jin2024hierarchical use Banzhaf interactions to enhance the training of video–language encoders. For a comprehensive overview of work in this direction, see [rozemberczki2022shapley, muschalik2024shapiq] and references given there.
@@ 3 A Game-theoretic Explanation of Similarity in Vision–Language Encoders
An interaction explanation can be viewed as a complete graph with weighted nodes and edges, similar to the SI-Graph [muschalik2025exact], or a symmetric quadratic matrix of dimension $n_{\mathcal{I}}+n_{\mathcal{T}}$, where the diagonal represents token attributions. To compute explanations, we mask tokens with a pre-defined baseline, depending on the encoder’s architecture, e.g. attention masking for text tokens (we defer implementation details to Appendix B). The masking operator is used to define the FIxLIP game that captures all possible masks.
@@ 3.1 FIxLIP-$\boldsymbol{p}$ explanations via weighted faithful Banzhaf interactions
In this section, we formally define the FIxLIP explanation as an instance of Definition˜1, which faithfully describes the similarity of all masked inputs $\nu(M)$. We thus define $\mathbf{e}$ as an additive explanation with interaction terms that recovers $\nu(M)\approx\hat{\nu}_{\mathbf{e}}(M)$ via a $2$-additive game [grabisch2016set] as
@@
The FIxLIP-$p$ explanation is the most faithful approximation of $\nu$ with respect to $\mathfrak{F}_{p}$. Its interactions correspond to the weighted Banzhaf interactions [marichal2011weighted]. For $p=0.5$, FIxLIP-$p$ is the faithful Banzhaf interaction index [tsai2023faith] with a known analytical solution [hammer1992approximations].
@@
Weighted Banzhaf interactions define a cardinal-probabilistic interaction index that satisfies the linearity, symmetry, and dummy axiom [fujimoto2006axiomatic]. Alternatively, faithful Shapley interactions optimize a variant of the faithfulness metric [tsai2023faith]. In our context, weighted Banzhaf interactions have two benefits: First and foremost, the hyperparameter $p$ allows for prioritizing masks that are expected to better reflect in-distribution inputs, and provides a flexible explanation framework with intuitive weights. Second, it is unknown if the Shapley weights can be factored into independent distributions for each modality – a crucial requirement for the cross-modal estimator introduced in Section˜3.2.
@@
For $p=0.5$, this estimator reduces to Faith-Banzhaf [tsai2023faith] – a variant of KernelSHAP [lundberg2017unified].
@@ 4 Evaluation Metrics for Interaction Explanations of VLE Predictions
In this section, we derive three evaluation metrics for explanations that may include second-order interactions. First, we evaluate the faithfulness of the approximation $\hat{\nu}_{\mathbf{e}}$ to $\nu$, since the basis of explaining VLEs is the masking operator that defines the FIxLIP game $\nu$. Note that conventionally this is done with an $R^{2}$ coefficient [kang2024learning, muschalik2024shapiq] or mean squared error [fumagalli2023shapiq, enouen2025instashap, musco2025provably], but we need to rely on rankings to compare attribution methods like saliency maps. Second, we generalize the area between insertion and deletion curves [AID, naofumi2023deletion, zhao2024gradeclip], aka remove least/most important first [bluecher2024decoupling], a popular faithfulness metric for token attribution methods. Finally, we extend the well-established pointing game evaluation of faithfulness from the image classification literature [bohle2021convolutional, bohle2024bcos, arya2024bcosification, moeller2025explaining]. In the latter two metrics, we need to assume $\mathbf{e}_{\{i,j\}}=0$ for the first-order attribution explanations without interactions.
@@ 5 Experiments
Setup. In experiments, we empirically validate the performance of FIxLIP with three metrics defined in Section 4, measure its computational efficiency, and demonstrate its utility in visual explanation of VLEs. We mainly use the openly available pre-trained CLIP models [radford2021learning] of two sizes: ViT-B/32 with $7\times 7$ image patches and ViT-B/16 with $14\times 14$. Moreover, we demonstrate the broader applicability of FIxLIP to explain SigLIP [zhai2023siglip] and SigLIP-2 [tschannen2025siglip2] up to the ViT-L/16 variant with $16\times 16$ patches. We rely on two openly available datasets commonly used in explainability research: MS COCO [lin2014microsoft] and ImageNet-1k [deng2009imagenet]; the latter specifically to design the pointing game evaluation considering zero-shot classification.
In quantitative evaluation, we compare FIxLIP to a few representative baselines: the first attribution method for CLIP abbreviated as GAME [chefer2021generic], a state-of-the-art first-order attribution method Grad-ECLIP [zhao2024gradeclip], and a recently released second-order interaction method exCLIP [moeller2025explaining]. For comparisons of Grad-ECLIP and exCLIP with previous baselines, refer to the appropriate benchmarks [zhao2024gradeclip, moeller2025explaining]. Figure 2 conceptually compares FIxLIP to baselines using the image–text input example from Figure 1. Furthermore, we naturally demonstrate the improvement between FIxLIP and first-order Shapley/Banzhaf values. In that, we effectively scale FIxLIP to approximate second-order explanations with a budget of over $10^{6}$ model inferences, which far exceeds related work. For FIxLIP-$p$, we use the cross-modal estimator with a budget of $2^{21}$, whereas FIxLIP with Shapley interactions uses the model-agnostic estimator with a budget of $2^{17}$, yielding approximately similar runtime. We mainly experiment with $p\in\{0.3,0.5,0.7\}$; note that there is no computational overhead for different $p$ settings. Further details for the experimental setup are provided in Appendix C.
Figure 2: Visual comparison between FIxLIP and baselines. First-order attribution methods, e.g. GAME and Grad-ECLIP, lack the tools to faithfully explain complex similarity predictions of vision–language encoders like CLIP. Notably, in this example, the text token antRefer to caption is the most important for the similarity prediction. One of the differences from exCLIP is that we include intra-modal and main effects in the approximation, which are crucial for obtaining faithful interaction explanations.